# RAFT: Adapting Language Model to Domain Specific RAG

### v1

@@ 4.2 Results
Using the above datasets and baselines, we evaluate our model RAFT  and demonstrate the effectiveness of RAFT  in Tab. 1. We see that RAFT  consistently and significantly outperforms the baselines. Compared with the base Llama-2 instruction-tuned model, RAFT  with RAG does much better in terms of extracting information as well as being robust towards distractors. The gain can be as big as 35.25% on Hotpot QA and 76.35% on Torch Hub evaluation. Compared with DSF on the specific dataset, our model does better at relying on the provided context to solve the problem. RAFT  does much better on tasks like HotpotQA and HuggingFace datasets (30.87% on HotpotQA and 31.41% on HuggingFace). Note that for PubMed QA, since it is a binary yes/no question, we don’t observe significant gains when we compare our model with DSF + RAG. Even compared with a much larger and better model GPT-3.5, RAFT  demonstrates significant advantages.
Overall, the LLaMA-7B model, both with and without the RAG, performs poorly due to its answering style not aligning with the ground truth. By applying domain specific tuning, we significantly enhance its performance. This process enables the model to learn and adopt the appropriate style of answering. However, introducing RAG to a domain-specifically fine-tuned (DSF) model doesn’t invariably lead to better outcomes. This might indicate that the model lacks training in context processing and extracting useful information from it. By incorporating our method, RAFT , we train the model not only to match its answering style with that required but also to improve its document processing capabilities. Consequently, our approach outperforms all others.
@@ 4.3 Effect of CoT
We also conduct an analysis to evaluate the effectiveness of the Chain-of-Thought approach in enhancing the model’s performance. As indicated in Table 2, simply providing the answer to a question may not always be adequate. This approach can lead to a rapid decrease in loss, resulting in the training process to diverge. Incorporating a reasoning chain that not only guides the model to the answer but also enriches the model’s understanding can improve the overall accuracy. In our experiments, integrating the Chain-of-Thought significantly enhances training robustness. We employ GPT-4-1106 to generate our Chain-of-Thought prompts and include an example of the prompt we used in Figure 3.
Table 2: Ablation on Chain-of-Thought: The numbers of RAFT  and RAFT  without CoT. Results on various datasets show that adding CoT can significantly improve the performance of the finetuned model. With a gain of 9.66% and 14.93% on the Hotpot QA and HuggingFace datasets respectively.
PubMed HotpotQA HuggingFace Torch Hub TensorFlow Hub
@@ 4.4 Qualitative Analysis
HotPot QA Question: What screenwriter with credits for “Evolution”[0/1879]e a film starring Nicolas Cage and Téa Leoni? Documents: $\ldots$ David Weissman is a screenwriter and director. His film credits include “The Family Man” (2000), “Evolution” (2001), and “When in Rome” (2010). $\ldots$ The Family Man is a 2000 American romantic comedy-drama film directed by Brett Ratner, written by David Diamond and David Weissman, and starring Nicolas Cage and Téa Leoni. … ——————————————————————————————————————————————– RAFT: ##Reason: The screenwriter with credits for the film “Evolution,” starring Nicolas Cage and Téa Leoni, is David Weissman. This information is provided in the reference documents which mention David Weissman as a screenwriter with film credits including “The Family Man” (2000), “Evolution” (2001), and “When in Rome” (2010). Therefore, the screenwriter for “Evolution” is David Weissman. ##Answer: David Weissman DSF: The Family Man
@@
Figure 4: Comparison of RAFT  and DSF: We prompt RAFT  and DSF fine-tuned models on the HotpotQA dataset. We can see that the DSF model extracts the wrong information from the context. For the question, who is the screenwriter, it responds with a film name. RAFT  manages to get the result correctly .
4.5 Should we train the LLM always with the oracle context for RAG?
@@
In our exploration of whether large language models (LLMs) should always be trained with the oracle context for Retrieval-Augmented Generation (RAG), we address a key question: what proportion (p%) of the training data should include oracle documents? Intuitively, one might assume that for effective training in reading and extracting information from context (e.g., RAG tasks), the oracle document should always be included during training (P = 100%). However, our findings challenge this assumption: incorporating a portion of the training data without the oracle document in the context (P = 80%) appears to enhance the model’s performance on RAG tasks.
@@
Fig. 5 presents our investigation into the hyperparameter P%, which represents the percentage of training instances that should include oracle documents. Our analysis reveals that the optimal proportion varies across datasets, with figures ranging from 40%, 60%, and 100%. This indicates that training your LLM without the correct corresponding context at times can be beneficial for the downstream task of answering questions related to the documents. In our training setup, we include four distractor documents alongside the oracle document, and at test time, we maintain this format by providing the oracle document with four distractors. Our findings suggest that, for domain specific RAG tasks, including a certain percentage of training data without the oracle documents in the context proves to be advantageous.
@@
Figure 5: How many golden documents to involve? We study the hyperparameter $P\%$ which indicates what fraction of the training data contains the oracle document(s) in its context. Results on NQ, TQA and HotpotQA suggest that mixing a fraction of data that does not have the oracle document in its context is helpful for in-domain RAG.
@@ 5 RAFT Generalizes to Top-K RAG
After demonstrating the performance of RAFT  on various benchmarks, we now study another important problem: How does the number of distractor documents in RAFT  affect the model’s performance when augmented with top-k retriever augmented generation (RAG) result during the evaluation? Previous research has highlighted the vulnerability of LLMs to irrelevant text (see studies  (Shi et al., 2023a; Weston & Sukhbaatar, 2023; Liu et al., 2023b)). This issue is particularly critical for LLMs + RAG since top-k RAG is frequently employed at test time to ensure high recall. Such a scenario necessitates the model to have the ability to discern and disregard irrelevant content, focusing solely on pertinent information.
@@ 5.1 Making Model Robust to top-K RAG
To tackle the challenge of enhancing large language models’ (LLMs) ability to sift through irrelevant text within the retrieval pipeline, our analysis revealed that training solely with oracle (highly relevant) documents can inadvertently diminish the model’s ability to discern and disregard irrelevant information. To address this, our algorithm, RAFT , adopts a strategy that integrates oracle documents with a mix of irrelevant ones. This methodology prompts us to investigate the ideal fraction of negative (irrelevant) documents to incorporate throughout the training process and to assess how well this training approach adapts to different volumes of documents encountered by the Retrieval-Augmented Generation (RAG) during the test phase. Our aim is to refine the balance between relevant and irrelevant information to strengthen the model’s efficiency in identifying and utilizing pertinent content. Notice that Sec 4.5 looked at what P% of training data should include distractors, while in this section, we study test-time scenarios.
Training with Negative Documents To enhance the robustness of large language models (LLMs) against irrelevant text in retrieved documents, we adopted a finetuning approach that incorporates both golden (highly relevant) documents and distractor (irrelevant) documents. The model was trained with varying numbers of distractor documents, but consistently evaluated using the top-k documents obtained from the retriever - not to be confused with $p$.
@@
Our findings, detailed in Fig. 6, reveal that finetuning with only the oracle document frequently results in inferior performance compared to configurations that include a greater number of distractor documents. As we can see in the figure, the better performance for Natural Questions is training with $D^{*}+3D$ and it is $D^{*}+1D$ documents with Hotpot QA. This insight has been particularly beneficial for our algorithm, RAFT . In our experiments, we typically employ a training setup consisting of one oracle document alongside four distractor documents. This approach strikes a balance, ensuring the model is not overwhelmed by distractors while still gaining the ability to effectively discern and prioritize relevant information.
Generalization to a variable number of test-time documents. We extended our research to examine the impact of different quantities of test-time documents on the model’s performance. Specifically, our experiments focused on assessing how models, trained with varying numbers of distractor documents, respond to changes in the number of documents presented at test time.
@@
The results, illustrated in Fig. 6, confirm that the inclusion of distractor documents during training indeed makes the model more resilient to fluctuations in the number of documents encountered during testing. This ability to maintain consistent performance despite variations in test-time document numbers further validates the robustness of our approach, RAFT . This finding underscores the importance of a well-calibrated training environment to prepare the model for a range of scenarios it may encounter in real-world applications.
@@
Figure 6: Test-Time Documents Varying: We study how robust RAFT is to varying numbers of test-time documents that a retriever might provide. In NQ, we find that training with 4 documents leads to the best performance, but training with 2 documents is optimal for HotpotQA. However, across both datasets, training with all datasets consisting of oracle documents hurts performance.
@@ 6 Related Works
Retrieval-Augmented Language Models RAG enhances language models by integrating a retrieval module that sources relevant information from external knowledge bases, significantly improving performance across various NLP tasks, including language modeling (Guu et al., 2020; Borgeaud et al., 2022; Khandelwal et al., 2019; Shi et al., 2023d; Lin et al., 2023b; Shi et al., 2023c; Asai et al., 2023; Xu et al., 2023; Wang et al., 2023) and open-domain question answering (Izacard et al., 2023; Lewis et al., 2020). This integration follows a “retrieve-and-read" paradigm where the retrieval module provides additional context from external sources, which the LM then uses to generate the final output. The retrieval process involves using the input as a query to fetch documents, which the LM incorporates for final predictions. For instance, Atlas (Izacard et al., 2023) fine-tunes T5 models with the retriever, treating documents as latent variables, while RETRO (Borgeaud et al., 2022) modifies the decoder-only architecture to include retrieved texts and conducts pre-training from scratch. kNN-LM (Khandelwal et al., 2019) interpolates between the LM’s next token distribution and distributions computed from retrieved tokens at inference. (Shi et al., 2023d; Ram et al., 2023) assume black-box access to an LM and combine it with either off-the-shelf or fine-tuned retriever.
Memorization A key question around large neural language models is whether they truly “understand” text  (Feldman, 2020; Power et al., 2022) or simply rely on surface pattern memorization  (Carlini et al., 2019; Tänzer et al., 2022). (Feldman, 2020; Carlini et al., 2019, 2022) develop methodologies to quantify the extent of memorization in neural models. (Brown et al., 2020; Power et al., 2022; Liu et al., 2022b) further explored how memorization impacts the models’ generalization capabilities. Recently, a seminal work by (Carlini et al., 2021; Shi et al., 2023b) demonstrated the ability of language models to memorize and regurgitate training data, raising significant privacy concerns (Kandpal et al., 2022; Pan et al., 2020).
Finetuning of LLMs Recent years have seen rapid progress in developing large-scale language models (LLMs) (Brown et al., 2020; OpenAI, 2023; Workshop et al., 2022; Touvron et al., 2023, 2023; Anil et al., 2023). To adapt these foundation models to downstream tasks, fine-tuning (Mishra et al., 2021; Sanh et al., 2021; Chung et al., 2022; Muennighoff et al., 2023; Zhou et al., 2023b; Lin et al., 2023b; Ji et al., 2024) has become a prevalent approach. Traditional supervised fine-tuning may be limited by the cost and compute required for adapating LLMs. Addressing these challenges, research in the realm of parameter-efficient fine-tuning (Houlsby et al., 2019), such as Prompt Tuning (Lester et al., 2021), Prefix-Tuning (Li & Liang, 2021), P-Tuning (Liu et al., 2022a) and Low-Rank based fine-tuning (Hu et al., 2021), has gained traction. These methods enable LLMs to acquire domain-specific knowledge and adapt to specialized tasks such as question answering, summarization, and dialogue generation. Another branch of finetuning is through RLHF (Ouyang et al., 2022; Rafailov et al., 2023; Liu et al., 2023a; Zhang et al., 2023), which adopts RL to align LLM’s preference with human.
@@ 7 Conclusion
RAFT is a training strategy designed to enhance the model’s performance in answering questions within a specific domain, in "open-book" settings. This technique demonstrates a fine-tuning recipe for LLMs for question-answering tasks based on a selected collection of documents. We have pinpointed several crucial design decisions, such as training the model alongside distractor documents, organizing the dataset so a portion lacks oracle documents in their context, and formulating answers in a chain-of-thought manner with direct quotations from the relevant text. Our evaluations on PubMed, HotpotQA, and Gorilla API Bench underline RAFT’s significant potential. Looking forward, we anticipate that in-domain Retrieval-Augmented Generation (RAG) will continue to gain interest within both industrial and academic spheres. Unlike general-RAG, our work addresses practical scenarios where LLMs are tasked with answering questions using domain-specific knowledge. Aligning with current trends, our findings suggest that smaller, fine-tuned models are capable of performing comparably well in domain-specific question-answering tasks, in contrast to their generic LLM counterparts.

### v2

@@ 4.1 Results
Using the above datasets and baselines, we evaluate our model RAFT  and demonstrate the effectiveness of RAFT  in Tab. 1. We see that RAFT  consistently and significantly outperforms the baselines. Compared with the base Llama-2 instruction-tuned model, RAFT  with RAG does much better in terms of extracting information as well as being robust towards distractors. The gain can be as big as 35.25% on Hotpot QA and 76.35% on Torch Hub evaluation. Compared with DSF on the specific dataset, our model does better at relying on the provided context to solve the problem. RAFT  does much better on the tasks like Hotpot and HuggingFace datasets (30.87% on Hotpot and 31.41% on HuggingFace). Note that for PubMed QA, since it is a binary yes/no question, we don’t observe significant gains when we compare our model with DSF + RAG. Even compared with a much larger and better model GPT-3.5, RAFT  demonstrates significant advantages.
Overall, the LLaMA-7B model, both with and without the RAG, performs poorly due to its answering style not aligning with the ground truth. By applying domain-specific tuning, we significantly enhance its performance. This process enables the model to learn and adopt the appropriate style of answering. However, introducing RAG to a domain-specifically fine-tuned (DSF) model doesn’t invariably lead to better outcomes. This might indicate that the model lacks training in context processing and extracting useful information from it. By incorporating our method, RAFT , we train the model not only to match its answering style with that required but also to improve its document processing capabilities. Consequently, our approach outperforms all others.
@@ 4.2 Effect of CoT
We also conduct an analysis to evaluate the effectiveness of the Chain-of-Thought approach in enhancing the model’s performance. As indicated in Table 2, simply providing the answer to a question may not always be adequate. This approach can lead to a rapid decrease in loss, resulting in the model beginning to overfit. Incorporating a reasoning chain that not only guides the model to the answer but also enriches the model’s understanding can improve the overall accuracy and prevent overfitting to concise answers. In our experiments, integrating the Chain-of-Thought significantly enhances training robustness. We employ GPT-4-1106 to generate our Chain-of-Thought prompts and include an example of the prompt we used in Figure 3.
Table 2: Ablation on Chain-of-Thought: The numbers of RAFT  and RAFT  without CoT. Results on various datasets show that adding CoT can significantly improve the performance of the finetuned model. With a gains of 9.66% and 14.93% in the Hotpot QA and HuggingFace datasets respectively.
PubMed HotpotQA HuggingFace Torch Hub TensorFlow
@@ 4.3 Qualitative Analysis
HotPot QA Question: What screenwriter with credits for “Evolution”[0/1879]e a film starring Nicolas Cage and Téa Leoni? Documents: $\ldots$ David Weissman is a screenwriter and director. His film credits include “The Family Man” (2000), “Evolution” (2001), and “When in Rome” (2010). $\ldots$ The Family Man is a 2000 American romantic comedy-drama film directed by Brett Ratner, written by David Diamond and David Weissman, and starring Nicolas Cage and Téa Leoni. … ——————————————————————————————————————————————– RAFT: ##Reason: The screenwriter with credits for the film “Evolution,” starring Nicolas Cage and Téa Leoni, is David Weissman. This information is provided in the reference documents where it mentions David Weissman as a screenwriter with film credits including “The Family Man” (2000), “Evolution” (2001), and “When in Rome” (2010). Therefore, the screenwriter for “Evolution” is David Weissman. ##Answer: David Weissman DSF: The Family Man
@@
Figure 4: Comparison of RAFT  and DSF: On the HotPot QA dataset, we can see that DSF model extracts the wrong information from the context when the question is asking who is the screen writer and it answers a film name. RAFT  manages to get the accurate results .
4.4 Should we train the LLM always with the golden context for RAG?
@@
In our exploration of whether large language models (LLMs) should always be trained with the golden context for Retrieval-Augmented Generation (RAG), we address a key question: what proportion (p%) of the training data should include golden documents? Intuitively, one might assume that for effective training in reading and extracting information from context (e.g., RAG tasks), the golden document should always be included during training (P = 100%). However, our findings challenge this assumption: incorporating a portion of the training data without the golden document in the context (P = 80%) appears to enhance the model’s performance on RAG tasks.
@@
Figure 5 presents our investigation into the hyperparameter P%, which represents the percentage of training instances that should include golden documents. We find that the optimal proportion varies across datasets, with P% ranging from 40%, 60%, and 100%. This indicates that training your LLM without the correct corresponding context at times can be beneficial for the downstream task of answering questions related to the documents. In our training setup, we include four distractor documents alongside the golden document, and at test time, we maintain this format by providing the golden document with four distractors. Our findings suggest that, for domain-specific RAG tasks, including a certain percentage of training data without the golden documents in the context proves to be advantageous.
@@
Figure 5: How many golden documents to involve? We study the hyperparameter P% where it indicates how much portion of training data is with golden document. Results on NQ, TQA and HotpotQA suggest that mixing some amount of data that the golden document is not put in the context is helpful for in-domain RAG.
@@ 5 RAFT Generalizes to Top-K RAG
We now study another important problem: How does the number of distractor documents in RAFT  affect the model’s performance when augmented with top-k RAG results during evaluation? Previous research has highlighted the vulnerability of LLMs to irrelevant text (see studies  (Shi et al., 2023a; Weston & Sukhbaatar, 2023; Liu et al., 2023)). This issue is particularly critical for LLMs + RAG since top-k RAG is frequently employed at test time to ensure high recall. Such a scenario necessitates the model to have the ability to discern and disregard irrelevant content, focusing solely on pertinent information.
@@ 5.1 Making Model Robust to top-K RAG
To tackle the challenge of enhancing large language models’ (LLMs) ability to sift through irrelevant text within the retrieval pipeline, our analysis revealed that training solely with golden (highly relevant) documents can inadvertently diminish the model’s ability to discern and disregard irrelevant information. To address this, our algorithm, RAFT , adopts a strategy that integrates golden documents with a mix of irrelevant ones. This methodology prompts us to investigate the ideal fraction of distractor (irrelevant) documents to incorporate throughout the training process and to assess how well this training approach adapts to different volumes of documents encountered by the Retrieval-Augmented Generation (RAG) during the test phase. Our aim is to refine the balance between relevant and irrelevant information to strenghten the model’s efficiency in identifying and utilizing pertinent content. Notice that Sec 4.4 looked what what P% of training data should include distractors, while in this section, we study test-time scenarios.
@@
Training with Distractor Documents To enhance the robustness of LLMs against irrelevant text in retrieved documents, we adopted a finetuning approach that incorporates both golden (highly relevant) documents and distractor (irrelevant) documents. The model was trained with varying numbers of distractor documents, but consistently evaluated using the top-3 documents obtained from the retriever - not to be confused with $p$. Our findings, detailed in Fig. 6, reveal that finetuning with only the golden document frequently results in inferior performance compared to configurations that include a greater number of distractor documents. As we can see in the figure, the better performance for Natural Questions is training with $D^{*}+3D$ and it is $D^{*}+1D$ documents with Hotpot QA. This insight has been particularly beneficial for our algorithm, RAFT . In our experiments, we consistently employ a training setup consisting of one golden document alongside four distractor documents.
@@
Generalization to a variable number of test-time documents. We extended our research to examine the impact of different quantities of test-time documents on the model’s performance. Specifically, our experiments focused on assessing how models, trained with varying numbers of distractor documents, respond to changes in the number of documents presented at test time. The results, illustrated in Fig. 6, confirm that the inclusion of distractor documents during training indeed makes the model more resilient to fluctuations in the number of documents encountered during testing. This ability to maintain consistent performance despite variations in test-time document numbers further validates the robustness of our approach, RAFT . This finding underscores the importance of a well-calibrated training environment to prepare the model for a range of scenarios it may encounter in real-world.
@@
Figure 6: Test-Time Documents Varying: To analyze how robust RAFT is to varying number of test-time documents, we study three domains – NQ, Trivia QA and HotPot QA. In NQ, we find that training with 4 documents leads to optimal performance, and this changes to 3 and 2 for for Trivia QA and HotPot QA respectively. However, we see that training with only golden documents leads to poor performance.
@@ 6 Related Works
Retrieval-Augmented Language Models Retrieval-Augmented Language Models (RALMs) enhance LLMs by integrating a retrieval module that sources relevant information from external knowledge bases, significantly improving performance across various NLP tasks, including language modeling (Guu et al., 2020; Borgeaud et al., 2022; Khandelwal et al., 2019; Shi et al., 2023d; Lin et al., 2023b; Shi et al., 2023c; Asai et al., 2023; Xu et al., 2023; Wang et al., 2023) and open-domain question answering (Izacard et al., 2023; Lewis et al., 2020). For instance, Atlas (Izacard et al., 2023) fine-tunes T5 models with the retriever, treating documents as latent variables, while RETRO (Borgeaud et al., 2022) modifies the decoder-only architecture to include retrieved texts and conducts pre-training from scratch. kNN-LM (Khandelwal et al., 2019) interpolates between the LM’s next token distribution and distributions computed from retrieved tokens at inference. (Shi et al., 2023d; Ram et al., 2023) assume black-box access to an LLM, combining it with either off-the-shelf or fine-tuned retriever.
Memorization A key question around large neural language models is whether they truly “understand” text  (Feldman, 2020; Power et al., 2022) or simply rely on surface pattern memorization  (Carlini et al., 2019; Tänzer et al., 2022). (Feldman, 2020; Carlini et al., 2019; 2022) develop methodologies to quantify the extent of memorization in neural models. (Brown et al., 2020; Power et al., 2022; Liu et al., 2022) further explored how memorization impacts the models’ generalization capabilities. (Carlini et al., 2021; Shi et al., 2023b) demonstrated the ability of language models to memorize and regurgitate training data, raising significant privacy concerns (Kandpal et al., 2022; Pan et al., 2020).
@@ 7 Conclusion
RAFT is a training strategy designed to enhance the model’s performance in answering questions within a specific domain, in "open-book" settings. We highlight several crucial design decisions, such as training the model alongside distractor documents, organizing the dataset so a portion lacks golden documents in their context, and formulating answers in a chain-of-thought manner with direct quotations from the relevant text. Our evaluations on PubMed, HotpotQA, and Gorilla API Bench underline RAFT’s significant potential.